DIRC Operations Weekly Report
Period: Wednesday 2007/2/14 - Wednesday 2007/2/21
Oncall: Jose Benitez
Fast Monitoring: Jose Benitez, Alejandro Perez
File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070221.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070221
PEP peak lumi: 6.5 x 1033
(previous week: ~7.1 x 1033)
BaBar integrated lumi: 1.7
(previous week: ~2.3 fb-1)
Run 6 PEP delivery:1.8
(previous week: ~2.4 fb-1)
Run 6 total delivery: 6.6 fb-1
Run 6 total BaBar recorded: 6.4 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
Foreseen work opportunities (PEP down periods):
BaBar activities/issues this week:
- Paged on Sunday evening due to transient GLINK error (5s) from sector 3. Problem could be related to DCC comunication with the ROMs.
logbook for details
- We have been paged on friday due to high backgrounds. Timing distributions in FM highly distorted.
- Biweekly Bkg Plot
- Replace two DFBs in sector 1 with innefficient or flaky channels (require bridge).
- HV group #11 in sector 0 is noisy: needs to be masked.
- Unmask all masked channels to find which of these are actually noisy. After December DFB work/movement masked channels may no longer be the noisy ones.
Bradley, Jose and Alejandro rotating backup & on-call commissioning duties